Mary is "y" years old. Her father is 4 times older than Mary. Her mother is 7 years younger than her father. If the three ages add up to 101 years, find the value of "y". Find the ages of Mary's parents.
Thanks in advance!

Found 2 solutions by elima, AnlytcPhil:
Answer by elima(1433)

Mary = y
Her father = 4y
Her mother = 4y-7
y+4y+4y-7=101
9y-7=101
9y=108
y=12
So Mary's age is 12, her father would be 4(12)=48,
Her maother would be 4(12)-7=48-7=41
Hope you understand
=)

Answer by AnlytcPhil(1806)

Mary is "y" years old. Her father is 4 times older than Mary. Her mother is 7
years younger than her father. If the three ages add up to 101 years, find the
value of "y". Find the ages of Mary's parents.
>>...Mary is "y" years old...<<
Translation: Mary's age = y
>>...Her father is 4 times older than Mary...<<
Translation: Father's age = 4y
>>...Her mother is 7 years younger than her father...<<
Transation: Mother's age = 4y - 7
>>...the three ages add up to 101 years...<<<
Translation: y + 4y + (4y - 7) = 101
Can you solve that for y? If not post again
Answers: y = 12, Mary is 12, her father is 48 and her mother is 41.
Edwin
AnlytcPhil@aol.com